Find the equation in slope-intercept form that passes through two points (1 , -1) and (-2, 2)
Answer by mananth(16946)


x1 y1 x2 y2
1 -1 -2 2

slope m = (y2-y1)/(x2-x1)
( 2 - -1 )/( -2 - 1
( 3 / -3 )
m= -1.00

Plug value of the slope and point ( 1 , -1 ) in
Y = m x + b
-1.00 = -1 + b
b= -1 - -1
b= 0
So the equation will be
Y = - x